OpenStudy (jojokiw3):

Can someone walk me through this exponential question and help check my individual steps?

OpenStudy (jojokiw3):

\[7^{3x-2} = 13^{x+1}\]

OpenStudy (jojokiw3):

Here's my first step: \[\log (7^{3x-2}) = \log(13^{x+1})\] What's should I try to do next?

OpenStudy (jojokiw3):

\[(3x-2) \log 7 = (x+1) \log 13\] Is that right?

OpenStudy (superdavesuper):

yup - next step is to expand the ( ) on both left n right; then re-group the two terms including x on 1 side n everything else on the other.

OpenStudy (jojokiw3):

Okay how exactly do I do it? Distributive property?

OpenStudy (superdavesuper):

correct :)

OpenStudy (jojokiw3):

question, can log 3x become 3 log x?

OpenStudy (jojokiw3):

That technically gets you the same answer right?

OpenStudy (superdavesuper):

nope: 3logx=log(x^3) in any case, plz note that the x is OUTSIDE of the log terms anyway

OpenStudy (jojokiw3):

Here's where I'm stuck, I don't know exactly how or why or how to use distributive here.

OpenStudy (superdavesuper):

i will help w the left hand side: 3xlog7 - 2log7 =.... plz do the right hand side

OpenStudy (jojokiw3):

log 7^3x - log 7^2?

OpenStudy (superdavesuper):

no need to put the 3x term back into the log function....leave it outside plz just do the same operation for the right hand side of the =......

OpenStudy (jojokiw3):

xlog13 + log13

OpenStudy (superdavesuper):

good :) so there are two terms; one each on the left and right that contains x. move them all on 1 side n the other two terms that do not contain x on the other side...

OpenStudy (jojokiw3):

3xlog 7 - xlog13 = 2log7 + log13

OpenStudy (superdavesuper):

VERY good :) now pull the common factor x out on the left side....

OpenStudy (jojokiw3):

\[x = \frac{ 2 \log 7 + \log 13 }{ 3 \log 7 - \log 13 }\]

OpenStudy (jojokiw3):

Right?

OpenStudy (superdavesuper):

bingo - u got it :)

OpenStudy (jojokiw3):

Ah thanks for the help. I think I got it now. :)

OpenStudy (superdavesuper):

welcome